Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Round your answer to four decimal places.)
μ = 4.3; σ = 1.8
P(3 ≤ x ≤ 6) =

Answer by stanbon(75887)

Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Round your answer to four decimal places.)
μ = 4.3; σ = 1.8
------
z(3) = (3-4.3)/1.8 = -0.7222
z(6) = (6-4.3)/1.8 = 0.9444
----------------------------
P(3 ≤ x ≤ 6) = P(-0.7222 < z < 0.9444) = 0.5924
---------------
Cheers,
Stan H.
----